Find the exact value solution
log (5x) + log (x-1) = 2

Answer by sachi(548)

log (5x) + log (x-1) = 2
or log[5x(x-1)]=2
or 5x(x-1)=AntiLog(2) = 102=100
or x(x-1)=100/5=20
or x^2-x-20=0
or x^2-5x+4x-20=0
or x(x-5)+4(x-5)=0
or (x-5)(x+4)=0
or x=5 or -4